OpenStudy (anonymous):

17.25 mL/s wat is it in cumecs (m^3/s)

OpenStudy (anonymous):

17.25ml = 0.000609178 cubic foot

OpenStudy (sirm3d):

use the conversion \(\displaystyle \large 1mL=1mm^3\)\[\Large 17.25\frac{\text{mL}}{s}=17.25\frac{\text{ mm}^3}{s}\times\left( \frac{1\text{ m}}{1000 \text{ mm}}\right)^3\]

OpenStudy (anonymous):

is it 0.0001725 cumecs

OpenStudy (anonymous):

One cumec is equivalent to 35.3147 cubic feet per second in the United States customary units.